Ammermann Components recently paid a dividend of $1 per share. This dividend is expected to grow at a rate of 25 percent a year for the next five years, after which it is expected to grow at a rate of 7 percent a year. The required rate of return for this stock is 18 percent. What is the estimated price of the stock?
Answer #1

D1=(1*1.25)=1.25

D2=(1.25*1.25)=1.5625

D3=(1.5625*1.25)=1.953125

D4=(1.953125*1.25)=2.44140625

D5=(2.44140625*1.25)=3.05175781

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(3.05175781*1.07)/(0.18-0.07)

=29.6852805

Hence price of stock=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.25/1.18+1.5625/1.18^2+1.953125/1.18^3+2.44140625/1.18^4+3.05175781/1.18^5+29.6852805/1.18^5

=$18.94(Approx).